# PLANesT-3D: A new annotated dataset for segmentation of 3D plant point clouds

### v1

@@ Abstract
Creation of new annotated public datasets is crucial in helping advances in 3D computer vision and machine learning meet their full potential for automatic interpretation of 3D plant models. In this paper, we introduce PLANesT-3D; a new annotated dataset of 3D color point clouds of plants. PLANesT-3D is composed of 34 point cloud models representing 34 real plants from three different plant species: Capsicum annuum, Rosa kordana, and Ribes rubrum. Both semantic labels in terms of "leaf" and "stem", and organ instance labels were manually annotated for the full point clouds. As an additional contribution, SP-LSCnet, a novel semantic segmentation method that is a combination of unsupervised superpoint extraction and a 3D point-based deep learning approach is introduced and evaluated on the new dataset. Two existing deep neural network architectures, PointNet++ and RoseSegNet were also tested on the point clouds of PLANesT-3D for semantic segmentation.
@@ 1 Introduction
Cultivation of horticultural crops is one of the most labor-intensive economic activities, and is central to global food security [1, 2]. The accuracy and efficiency of horticultural practices, such as pruning, weed control, disease management, and harvesting directly influence the yield and quality of the crops [3]. The increasing global demand for high-quality crops, hence the demand for intensive labor motivates development of technologies for automation in horticulture [4, 5]. Automated horticultural practices are expected to adapt to plant traits, which are diverse across plant species [6, 7]. Automation of trait measurements is also key to high throughput plant phenotyping, a fundamental process in plant research, breeding, and crop management [8, 9, 10]. Modeling of traits within a population, and capturing the structure of single plants or plant communities for particular operations have become major applications of computer vision [11, 12, 13, 14].
3D computer vision is particularly relevant in many automated horticultural operations. Robotic systems designed to perform pruning, weeding, and harvesting require correct identification of plant parts, and accurate estimation of the location, size, shape, and orientation of the structures in 3D space [15, 16, 17, 18]. Advances in 3D sensing technologies and machine learning fuel the development of algorithms for automatic segmentation of plants into their structural parts as well as organ-level trait estimation [19, 20].
Despite the considerable progress in 3D geometry modeling and analysis of plants in the last decades [21, 22], creation of benchmark datasets, a fundamental component for advancement of relevant research, is still lagging behind compared to other domains [23, 24]. Annotated public datasets of plant models are of particular importance both for comparison of competing algorithms and also for providing training data to machine learning processes [13].
@@
[25] Maize Instance
@@
 [27] Tobacco Growth
[28] Sorghum conditions
[29] Arabidopsis
@@
In the last decade, open datasets of annotated 2D plant images have been proliferated to assist evaluation of performance and development of machine learning models for plant phenotyping [31]. However, open datasets containing annotated and complete 3D plant models are rare. Table 1 gives a list of such existing datasets. Plant3D [27, 28, 29] contains a total of 714 3D laser scans of Tomato, Tobacco, Sorghum, and Arabidopsis plants obtained within 20-30 days of development. Information on the growth conditions of the plants are also provided. The 3D models do not contain color information. [32] manually labeled 54 models of the Plant3D data into stem and lamina points, and evaluated a range of machine learning approaches on this subset. [33] annotated 546 of the plants of the Plant3D dataset with semantic and instance labels. The labels are available upon request from [33]. ROSE-X dataset [26] contains 3D models of 11 rosebush plants acquired through X-ray computer tomography. The voxels in the volumetric models were labeled into three semantic categories: "Leaf", "Stem", and "Flower". The point cloud versions of the volumetric models are also available. Pheno4D [25] is a dataset of 3D point clouds of 7 maize plants and 7 tomato plants. The plants were scanned with a laser scanner at different growth stages resulting in 244 point clouds. 126 of them were manually annotated with semantic and instance labels. The Soybean-MVS dataset [30] is fundamentally different from the rest in terms of data acquisition modality. The plants were captured with an RGB camera in a controlled setup and their corresponding point clouds were created through multi-view stereo. A total of 102 point cloud models of five different soybean varieties were reconstructed at 13 stages of the whole growth period.
@@
These 3D plant datasets have been often used for assessing off-the-shelf machine learning tools for semantic and instance segmentation [32, 25, 34, 30]. The datasets also enabled training and evaluation of new approaches specifically developed for plant data [35, 36, 37, 38], including deep learning architectures such as PSegNet [33], PlantNet [39], RoseSegNet [40], and FF-Net [41].
@@
Although smaller in terms of the number of models as compared to Pheno4D, Plant3D, and Soybean-MVS (Table 1), the new dataset (PLANesT-3D  [^1] The PLANesT-3D dataset is publicly available at https://www.dropbox.com/s/8gl0r0zo0j6kmdo/PLANesT-3D.rar?dl=0) introduced in this work contributes to the diversity of publicly available annotated 3D plant models in terms of acquisition modality, number of distinct plants, and inclusion of other species. The data is acquired with an RGB camera and reconstructed through structure from motion (SfM) and multi-view steoroscopy (MVS). Together with the Soybean-MVS dataset, the PLANesT-3D dataset will balance the dominant acquisition modality, which is laser scanning. Reconstruction with SfM and MVS does not require expensive equipment or elaborate acquisition setups, while yielding both geometric and textural information.
@@
The objective of this publication is twofold: First is to introduce PLANesT-3D, a new dataset containing complete 3D color point clouds of plants together with their semantic and instance labels. Second is to present a new segmentation approach, the SP-LSCnet, that combines an unsupervised clustering scheme and an adaptive network for point cloud classification. We also provide semantic segmentation results on the dataset with two point-based deep learning architectures, PointNet++ and RoseSegNet. The main contributions of this work can be summarized as follows:
@@
    We propose SP-LSCnet as a new semantic segmentation approach where an unsupervised clustering method based on t-distributed stochastic embedding (t-SNE) is combined with a point cloud classifier network. The classifier network employs two adaptive modules to adjust local region organization for feature extraction.
@@
    We tested RoseSegNet, a semantic segmentation algorithm we had previously developed for plant data [40] on the new dataset, and demonstrated that it is effective without requiring hyperparameter re-adjustment.
@@ 2.1 Data acquisition
In order to obtain full 3D plant models, 2D color images of 34 plants were acquired using a handheld DSLR camera (EOS 6D MARK II, Canon, Japan). The acquisition was performed on 10 potted pepper plants (Capsicum annuum), 10 potted rose plants (Rosa kordana), and 14 ribes plants (Ribes rubrum), planted directly to the soil in a greenhouse. Multiple color images of resolution $6240\times 4160$ were captured manually around each plant via positioning the camera so as to cover as much plant surface as possible with a high degree of overlap. Three examples of the sets of camera locations and orientations are illustrated in Fig. 1.
Table 2 gives the number of images used to reconstruct the 3D plant point clouds. Three pairs of Ribes were very closely planted (Ribes 02 & 12, Ribes 04 & 13, Ribes 05 & 14). A single point cloud was reconstructed for each pair using the set of images where the pair appeared together. After reconstruction, noise removal and pose normalization, the two plants were manually segmented into separate point clouds of two individual plants.
@@ 2.2 Reconstruction and pre-processing of 3D point clouds
For reconstruction of 3D color point clouds from 2D color images, Agisoft Metashape Professional (Agisoft LLC, St. Petersburg, Russia) was employed. Agisoft Metashape is a software that performs photogrammetric processing of digital images and generates 3D models through principles of structure from motion (SFM) and multi-view reconstruction [42]. Examples of the raw point clouds produced by Metashape Professional are given in Fig. 2. The software also provides confidence maps for the 3D points. The confidence value for a point corresponds to the number of images that "see" and contribute to reconstruct the point.
@@
In order to recover the correct scale of the point cloud, remove background structures and noise, and segment the plant points from the scene, we devised a simple, semi-automatic procedure. We denote the raw point cloud as $\mathcal{P}_{R}$, where a point $p\in\mathcal{P}_{R}$ is represented by its coordinates $p=(x,y,z)$ defined in the coordinate frame returned by Metashape. The steps of the procedure are detailed as follows:
[…]
@@
  • Rotating the point cloud to a normalized pose: We aimed to rotate the reference frame such that the XY-plane corresponded to the ground or the plane holding the plant pot. In order to estimate the parameters of this plane, we used M-estimator SAmple Consensus (MSAC) algorithm given in [43], which is a variant of RANdom SAmple Consensus (RANSAC) algorithm. We manually picked a point, $p^{\prime}_{base}$, close to the plant base. Using the estimated normal vector $\mathbf{n}=[n_{x},n_{y},n_{z}]$ of the plane in the scene and the point $p^{\prime}_{base}$ at the plant base, we translated and rotated the point cloud $\mathcal{P}^{\prime}_{R}$ into a new reference frame using
@@
    for all $p^{\prime}\in\mathcal{P}^{\prime}_{R}$. The point cloud rotated to this normalized pose is denoted as $\hat{\mathcal{P}}_{R}$. In Eq. 3, $\mathbf{R}$ corresponds to the rotation matrix which transforms the normal $\mathbf{n}$ to $[0,0,1]$ in the new reference frame. The origin of this reference frame corresponds to the hand-picked point at the plant base. The XY plane coincides with the plane detected by the MSAC algorithm. The positive Z-direction is oriented from the plant base towards the shoot.
  • Extracting plant points: Once the origin of the reference coordinate frame is established at the plant base, and the positive Z-direction points toward the plant shoot, we removed all points with a negative $z$ coordinate, assuming all plant points remain above the plant base. We verified this assumption for all the models in the dataset. We applied connected component analysis to the rest of the point cloud and retained the points belonging to the largest connected component as the final plant point set $\mathcal{P}\subset\hat{\mathcal{P}}_{R}$. Examples of the outputs of this semi-automatic process is given in Fig. 2-(b) for a rose plant and in Fig. 2-(d) for a ribes plant.

### v2

@@ Abstract
Creation of new annotated public datasets is crucial in helping advances in 3D computer vision and machine learning meet their full potential for automatic interpretation of 3D plant models. Despite the proliferation of deep neural network architectures for segmentation and phenotyping of 3D plant models in the last decade, the amount of data, and diversity in terms of species and data acquisition modalities are far from sufficient for evaluation of such tools for their generalization ability. To contribute to closing this gap, we introduce PLANesT-3D; a new annotated dataset of 3D color point clouds of plants. PLANesT-3D is composed of 34 point cloud models representing 34 real plants from three different plant species: Capsicum annuum, Rosa kordana, and Ribes rubrum. Both semantic labels in terms of "leaf" and "stem", and organ instance labels were manually annotated for the full point clouds. PLANesT-3D introduces diversity to existing datasets by adding point clouds of two new species and providing 3D data acquired with the low-cost SfM/MVS technique as opposed to laser scanning or expensive setups. Point clouds reconstructed with SfM/MVS modality exhibit challenges such as missing data, variable density, and illumination variations. As an additional contribution, SP-LSCnet, a novel semantic segmentation method that is a combination of unsupervised superpoint extraction and a 3D point-based deep learning approach is introduced and evaluated on the new dataset. The advantages of SP-LSCnet over other deep learning methods are its modular structure and increased interpretability. Two existing deep neural network architectures, PointNet++ and RoseSegNet, were also tested on the point clouds of PLANesT-3D for semantic segmentation.
@@ 1 Introduction
The cultivation of horticultural crops is one of the most labor intensive economic activities and is central to global food security [Riisgaard2011, van2016horticultural]. The accuracy and efficiency of horticultural practices, such as pruning, weed control, disease management, and harvesting, directly influence the yield and quality of crops [tyagi2017pre]. The increasing global demand for high-quality crops, and hence the demand for intensive labor, motivates the development of automation technologies in horticulture [Bac2014, bogue2016robots]. Automated horticultural practices are expected to adapt to plant traits, which are diverse across plant species [ZHANG2023, vanHenten2022]. Automation of trait measurements is also key to high throughput plant phenotyping, a fundamental process in plant research, breeding, and crop management [atefi2021, Pieruschka2019, James2022]. Modeling of traits within a population and capturing the structure of individual plants or plant communities for particular operations have become major applications of computer vision [Minervini2015, KOLHAR2023, Coleman2023, Zieschank2023community].
3D computer vision is particularly relevant in many automated horticultural operations. Robotic systems designed to perform pruning, weeding, and harvesting require correct identification of plant parts and accurate estimation of the location, size, shape, and orientation of the structures in 3D space [kootstra2021, RONG2022, ZAHID2021, LELOUEDEC2021]. Advances in 3D sensing technologies and machine learning fuel the development of algorithms for automatic segmentation of plants into their structural parts, as well as organ-level trait estimation [gibbs2016approaches, paulus2019measuring].
Despite the considerable progress in 3D geometry modeling and analysis of plants in the last decades [Okura2022, harandi2023make], creation of benchmark datasets and provision of publicly available annotated data for training machine learning techniques, which are fundamental components for advancement of relevant research, are still lagging behind compared to other domains [BARTH2018, Chaudhury2020]. Annotated public datasets of plant models are of particular importance both for comparison of competing algorithms and also for providing training data to machine learning processes [Coleman2023].
@@
[Schunck2021pheno4d] Maize Instance
@@
[Conn2017high] Tobacco Growth
[Conn2017statistical] Sorghum conditions
[Conn2019] Arabidopsis
@@
In the last decade, open datasets of annotated 2D plant images have been proliferated to assist evaluation of performance and development of machine learning models for plant phenotyping [LU2020dataSurvey]. However, open datasets containing annotated and complete 3D plant models are rare. Table 1 gives a list of such datasets. Plant3D [Conn2017high, Conn2017statistical, Conn2019] contains a total of 714 3D laser scans of Tomato, Tobacco, Sorghum, and Arabidopsis plants obtained within 20-30 days of development. Information on the growth conditions of the plants is also provided. The 3D models do not contain color information. [Ziamtsov2019] manually labeled 54 models of the Plant3D data into stem and lamina points, and evaluated a range of machine learning approaches on this subset. [Li2022PSegNet] annotated 546 of the plants of the Plant3D dataset with semantic and instance labels. The ROSE-X data set [Dutagaci2020rose] contains 3D models of 11 rosebush plants acquired by X-ray computer tomography. The voxels in the volumetric models were labeled into three semantic categories: "Leaf", "Stem", and "Flower". The point cloud versions of the volumetric models are also available. Pheno4D [Schunck2021pheno4d] is a dataset of 3D point clouds of 7 maize plants and 7 tomato plants. The plants were scanned with a laser scanner at different stages of growth, resulting in 244 point clouds. 126 of them were manually annotated with semantic and instance labels. The Soybean-MVS dataset [Sun2023-Soybean] is fundamentally different from the rest in terms of data acquisition modality. The plants were captured with an RGB camera in a controlled setup, and their corresponding point clouds were created through multi-view stereo. A total of 102 point cloud models of five different soybean varieties were reconstructed at 13 stages of the whole growth period.
@@
These 3D plant datasets have been often used for assessing off-the-shelf machine learning tools for semantic and instance segmentation [Ziamtsov2019, Schunck2021pheno4d, Turgut2022segmentation, Sun2023-Soybean]. The datasets also enabled training and evaluation of new approaches specifically developed for plant data [Ziamtsov2020, mirande2022, Xiang2022research, dutagaci2023using], including deep learning architectures such as PSegNet [Li2022PSegNet], PlantNet [Li2022plantnet], RoseSegNet [Turgut2022rosesegnet], and FF-Net [Guo2023FFnet].
@@
Although smaller in terms of the number of models as compared to Pheno4D, Plant3D, and Soybean-MVS (Table 1), the new dataset (PLANesT-3D  [^1] The PLANesT-3D dataset is publicly available at https://aperta.ulakbim.gov.tr/record/286354 and https://github.com/visionlab-ogu/PLANesT-3D/tree/main/data) introduced in this work contributes to the diversity of publicly available annotated 3D plant models in terms of acquisition modality, number of distinct plants, and inclusion of other species. The data is acquired with an RGB camera and reconstructed through structure from motion (SfM) and multi-view steoroscopy (MVS). Together with the Soybean-MVS dataset, the PLANesT-3D dataset will balance the dominant acquisition modality, which is laser scanning. Reconstruction with SfM and MVS does not require expensive equipment or elaborate acquisition setups, while yielding both geometric and textural information.
@@
The objective of this study is twofold: First is to introduce PLANesT-3D, a new dataset containing complete 3D color point clouds of plants together with their semantic and instance labels. Second is to present a new segmentation approach, the SP-LSCnet [^2] The code for SP-LSCnet is available at https://github.com/visionlab-ogu/PLANesT-3D, that combines an unsupervised clustering scheme and an adaptive network for point cloud classification.
As the capacity of machine learning techniques to model a large range of variations in data increases rapidly, more and more training samples are required to exploit the potential of those techniques. In the domain of 3D plant phenotyping, both the diversity and amount of labeled data acquired from real plants are very important. Even for a single plant species, more data is required for machine learning techniques to model the variations among plant instances in terms of structural and geometric complexity. PLANesT-3D introduces new instances from three different species, with a large within-species variety in terms of plant size, architectural complexity, and the number and density of organs, such as leaves.
Introducing annotated 3D models of different species is also key to train machine learning techniques and compare their merits and shortcomings over multiple species. In this respect, PLANesT-3D contributes to the field with annotated models of two new species, pepper and ribes. These species, especially ribes, bring in more variability in terms of size, architecture, and geometric variations of the organs. The ribes plants in our dataset vary in size from 30 cm to 1m. The leaves are many in quantity (reaching 100 for some individual plants) and dense compared to plants in existing datasets. They also possess distinct geometric structures. The diversity in terms of the shapes of the leaves and stems of pepper, rose and ribes plants enriches the collection of available datasets.
Provision of annotated 3D models of various species is indispensable to assess the generalization ability and transferability of a machine learning model to instances of a different species. Also, such trained models can be re-trained for comparison of various domain-adaptation techniques with respect to the amount of new annotated data they require to effectively process instances of a new species. Although not explored in this study experimentally, we envision that PLANesT-3D will enable researchers to test their machine learning models on different species, instances of which were not present in their training sets.
Plant phenotyping and plant monitoring systems operate with acquisition devices ranging from single cameras and multi-camera setups to RGB-D sensors and laser scanners. The nature of the acquired 3D data differs widely depending on the sensor in terms of point density, and the structure and amount of noise. While simulating noise and varying point density is possible by virtual manipulation of existing point clouds, 3D models originating from various acquisition systems are precious to assess the robustness of phenotyping systems. As distinct from existing datasest, PLANesT-3D dataset introduces 3D plant models built from images acquired by a single hand-held camera. This particular reconstruction technique resulted in point clouds which contain highly-dense and high-quality regions together with sparse and noisy regions due to heavy self-occlusion.
The second objective of this study is to introduce a new semantic segmentation approach, where the 3D data points are mapped to 2D domain for clustering and the corresponding 3D point cloud of each cluster is processed separately with a classification network. State-of-the-art deep learning methods processing plant point clouds are mostly designed as end-to-end networks, whose layers and modules are tuned to the training data. Also, such systems operate as a black-box and do not offer much interpretability to the end-user.
We would like to showcase an alternative to this approach, the SuperPoint-Leaf-Stem-Classification network (SP-LSCnet), where a portion of the system is unsupervised, easy to visualize, and interpretable. The unsupervised clustering scheme maps the 3D point cloud into the 2D space via t-distributed stochastic neighborhood embedding (t-SNE), which clusters points on the same manifold together and preserves the local structure of the clusters. This behavior of t-SNE is well-suited for 3D plant data, which, when mapped to 2D, looks like flattened versions of parts of 3D leaves and branches. The subsequent unsupervised clustering method allows the user to visualize the clusters in a single map, where the nature of each cluster is visible. The method does not necessitate training, and is readily applicable to new plant species.
The subsequent classification stage in the new SP-LSCnet method operates on the 3D counterparts of the clusters obtained in 2D. Each cluster is processed as a separate point cloud and processed by an adaptive network for classification into a leaf or a stem. The advantage of this modular approach is that the clusters can be processed by any other classification network, allowing the user to track and compare the results visually on the 2D map of clusters. It is also possible to convert this visualization ability into an interface where the user can correct wrong predictions. Such an interface can be a basis of a fast annotation tool, where the annotator can observe a complex 3D plant point cloud through a 2D mapping of organs and parts of organs with class labels predicted by a network and can correct classification errors quickly.
In summary, PLANesT-3D is provided to help close the gap between data scarcity in terms of amount and diversity of 3D plant models and the progress on the machine learning methods for segmentation of 3D plant models. PLANesT-3D introduces annotated point clouds of plants of two new species, which were not present in the existing datasets. Also, PLANesT-3D, as opposed to the existing datasets, provides point clouds that were acquired with a low-cost modality that do not require specialized setups. In this way, PLANesT-3D will contribute to the diversity of available data for assessing generalization ability of machine learning methods for 3D plant phenotyping. We also introduce the SP-LSCnet method to demonstrate that a partly-unsupervised segmentation scheme, with intermediate 2D visualization options yields competitive performance. These intermediate 2D visualizations greatly increase the interpretability of our method in contrast to black-box deep neural networks. Its unsupervised nature and modularity increases its ability to generalize to new species without extensive training. In addition, we provide semantic segmentation results on the PLANesT-3D dataset with two point-based deep learning architectures, PointNet++ and RoseSegNet.
The main contributions of this work can be summarized as follows:
@@
    We propose SP-LSCnet as a new semantic segmentation approach where an unsupervised clustering method based on t-SNE is combined with a point cloud classifier network. The classifier network employs two adaptive modules to adjust local region organization for feature extraction.
@@
    We tested RoseSegNet, a semantic segmentation algorithm we had previously developed for plant data [Turgut2022rosesegnet] on the new dataset, and demonstrated that it is effective without requiring hyperparameter re-adjustment.
@@ 2.1 Data acquisition
Plants have intricate architectures and individuals of a single species possess significant diversities in terms of structure, shape and texture. The number of individual plants and the representativeness of the variations in a training dataset are essential for machine learning algorithms to model a specific species effectively. However, generating manually annotated data on dense and complicated point clouds is time intensive and the number of models to be annotated is restricted by available human resources.
We targeted to reconstruct and annotate 10 individual plants for each of the three species for our new dataset: pepper (Capsicum annuum), rose (Rosa kordana), and ribes (Ribes rubrum), which are considered to be valuable produce. Rose models were already available in the ROSE-X dataset [Dutagaci2020rose], however they were acquired with X-ray imaging. The new rose models in our dataset possess reconstruction noise and variations of point density. Models of the other two species were not present in other datasets. While selecting individual plants, we considered variations in terms of plant size, general shape and architecture, number of leaves, and the shape and texture of the branches and leaves. As a result of this selection process, we modeled 10 potted pepper plants, 10 potted rose plants, and 14 ribes plants planted directly to the soil in a greenhouse.
In order to obtain full 3D plant models, 2D color images of 34 plants were acquired using a handheld DSLR camera (EOS 6D MARK II, Canon, Japan). Multiple color images of resolution $6240\times 4160$ were captured manually around each plant via positioning the camera so as to cover as much plant surface as possible with a high degree of overlap. Three examples of the sets of camera locations and orientations are illustrated in Fig. 1.
Table 2 gives the number of images used to reconstruct the 3D plant point clouds. Three pairs of Ribes were very closely planted (Ribes 02 & 12, Ribes 04 & 13, Ribes 05 & 14). A single point cloud was reconstructed for each pair using the set of images where the pair appeared together. After reconstruction, and pose normalization, the two plants were manually segmented into separate point clouds of two individual plants.
@@ 2.2 Reconstruction and pre-processing of 3D point clouds
For reconstruction of 3D color point clouds from 2D color images, Agisoft Metashape Professional (Agisoft LLC, St. Petersburg, Russia) was employed. Agisoft Metashape is a software that performs photogrammetric processing of digital images and generates 3D models through principles of structure from motion (SFM) and Multi-view Stereo (MVS) [hartley2003multiple]. Examples of the raw point clouds produced by Metashape Professional are given in Fig. 2. The software also provides confidence maps for the 3D points. The confidence value for a point corresponds to the number of images that "see" and contribute to reconstruct the point.
@@
In order to recover the correct scale of the point cloud, remove background structures and noise, and segment the plant points from the scene, we devised a simple, semi-automatic procedure. The manual operations involve the following: 1) Picking landmark points from the raw 3D point cloud to recover the scale. 2) Measuring the distances between landmark points in the real scene. 3) Picking a point on the plant close to the plant base. The first two operations are used to scale the point cloud with the scale factor, which is the ratio of the measurements on the point cloud and on the real scene. The recovery of scale through landmark points is a common practice for 3D reconstruction with SfM. The third operation is used to mark the center of the new coordinate frame. Apart from these three manual interventions, the rest of the procedure is performed automatically.
We denote the raw point cloud as $\mathcal{P}_{R}$, where a point $p\in\mathcal{P}_{R}$ is represented by its coordinates $p=(x,y,z)$ defined in the coordinate frame returned by Metashape. The steps of the procedure are detailed as follows:
  • Recovering the scale: The multi-view reconstruction pipeline employed by Metashape Professional provides the scene geometry up to a scale, meaning that the correct scale of the structure is lost. To enable recovery of the correct scale, we installed rulers and/or objects with distinct patterns in the vicinity of the plants during image acquisition. Once the 3D point cloud is reconstructed, a number of landmarks such as the markers on the rulers and/or on the patterns were picked manually. Let the set of these landmark points be $\{p_{r}\}\subset\mathcal{P}_{R}$ with $r=1,2,...,N_{r}$, where $N_{r}$ is the number of landmark points. The distance between a pair of these reference points on the point cloud was determined as $\hat{d}_{rs}=\|p_{r}-p_{s}\|$, where $\|\cdot\|$ corresponds to the Euclidean norm. The average of the distances of all the pairs is equal to
@@
  • Rotating the point cloud to a normalized pose: We aimed to rotate the reference frame such that the XY-plane corresponded to the ground or the plane holding the plant pot. In order to estimate the parameters of this plane, we used M-estimator SAmple Consensus (MSAC) algorithm given in [Torr2000], which is a variant of RANdom SAmple Consensus (RANSAC) algorithm. We manually picked a point, $p^{\prime}_{base}$, close to the plant base. Using the estimated normal vector $\mathbf{n}=[n_{x},n_{y},n_{z}]$ of the plane in the scene and the point $p^{\prime}_{base}$ at the plant base, we translated and rotated the point cloud $\mathcal{P}^{\prime}_{R}$ into a new reference frame using
@@
    for all $p^{\prime}\in\mathcal{P}^{\prime}_{R}$. The point cloud rotated to this normalized pose is denoted as $\hat{\mathcal{P}}_{R}$. In Eq. 3, $\mathbf{R}$ corresponds to the rotation matrix which transforms the normal $\mathbf{n}$ to $[0,0,1]$ in the new reference frame. The origin of this reference frame corresponds to the hand-picked point at the plant base. The XY plane coincides with the plane detected by the MSAC algorithm. The positive Z-direction is oriented from the plant base towards to the shoot.
  • Extracting plant points: Once the origin of the reference coordinate frame is established at the plant base, and the positive Z-direction points toward the plant shoot, we removed all points with a negative $z$ coordinate, assuming all plant points remain above the plant base. We verified this assumption for all the models in the dataset. We applied connected component analysis to the rest of the point cloud and retained the points belonging to the largest connected component as the final plant point set $\mathcal{P}\subset\hat{\mathcal{P}}_{R}$. Examples of the outputs of this semi-automatic process is given in Fig. 2b for a rose plant and in Fig. 2d for a ribes plant.
In summary, the automatic operations in this semi-automatic method involve 1) detecting a plane through MSAC to establish the orientation of the XY plane of the reference frame; 2) using the manually picked point close to the plant base to direct the positive direction of Z axis towards the plane shoot, and translate the origin of the reference frame; 3) removing points below the XY plane; and 4) applying connected components analysis to select the largest component as the plant point cloud. One assumption for this procedure to work for other settings of 3D plant data acquisition is the availability of a prominent plane (the ground or a platform) that is relatively perpendicular to the plant orientation. Another assumption is that the plant parts should not trail below the supporting surface.
Figure 3: Sample point clouds from PLANesT-3D demonstrating the diversity in terms of the overall plant shape and architecture (a). Examples of dense foliage in the PLANesT-3D dataset (b).
Figure 4: Diversity in leaf shape and texture (a). Diversity in stem shape (b).
Figure 5: Instances of noisy and missing data due to self-occlusion and reconstruction errors.
2.3 The PLANesT-3D dataset
The PLANesT-3D dataset consists of 34 plant point clouds, where the plant points are isolated from the background as described in the previous section. Instances of three species are present in the dataset. The global shape and architecture, especially for the Rose and Ribes plants are diverse in PLANesT-3D as demonstrated in Fig. 3a. The number of leaves vary between 21 and 105, and the height of the plants is in the relatively large range of 25 to 105cm. The density of the foliage varies from plant to plant, with some plants having high-density foliage as shown in Fig. 3b. The closely packed leaves pose a particular challenge for leaf instance segmentation methods.
The PLANesT-3D dataset is rich in the geometric and textural variations of leaf and stem instances. Fig. 4a shows isolated leaves of high point density from the dataset. The leaves of Ribes plants are of considerable diversity, both in shape and appearance. Stems of Ribes and Rose plants have diverse shape characteristics as illustrated in Fig. 4b, such as the presence of cut stems, varying stem thickness and curvature, varying branching angles, and variety in the arrangements of thorns.
We selected high point density leaves for illustrating the shape variations in Fig. 4a. However, due to the presence of thin leaves and branches and heavy self-occlusion around high density foliage, the point clouds have structures with missing data and reconstruction noise. Instances of such cases are given in Fig. 5. Noise is especially present at boundaries with color bleeding from the background. Color bleeding is the phenomena where colors from one area of the scene spreads into other objects, especially into the borders. There is also reconstruction noise present on the surfaces of leaves and stems. Parts of heavily occluded branches and leaves were not observed from any of the camera positions, leading to missing portions in the 3D models. Some small leaves and thin branches were not fully reconstructed resulting from the limited resolution of the 2D images.